Life Insurance A 39-year-old woman purchases a $300,000 term life insurance policy for an annual payment of $450. Based on a period life table for the U.S. government, the probability that she will survive the year is 0.999053Find the expected value of the policy for the insurance companyRound to two decimal places for currency problems


1
Expert's answer
2022-03-29T07:27:44-0400

"EX=\left( 450-300000 \right) \cdot \left( 1-0.999053 \right) +450\cdot 0.999053=165.9"